OpenStudy (anonymous):

Part 2: Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs (5, 1), (–3, 4) (0, –2), (3, 2) (–2, –1), (1

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

nvm thas ok i got the answer it's : (5,1) and (-3,4) slope = (4-1)/(-3-5) = 3/-8 or -3/8 using (5,1) y-1 = (-3/8)(x-5) from y - y1 = m(x - x1) at this point multiply each side by the denominator of the slope, if the slope is a fraction 8y - 8 = -3(x-5) your fractions have disappeared, yeahhh! 8y - 8 = -3x + 15 3x + 8y = 23 at this stage I use the point that was not used and test if it satisfies my equation. for (-3,4) LS = 3(-3) + 8(4) = -9+32 = 23 RS = 23 , all is good! Once you have the equation is the simple form of 3x + 8y = 23, you can go to any of the others changing it to slope - y intercept form takes 2 steps 1. keep only the y term on the left side 3x + 8y = 23 -------> 8y = -3x + 23 2. divide each term by the coeffiecient of the y term ------> y = (-3/8)x) + 23/8